OpenStudy (anonymous):

Solve the system of equations. 2x-4y+3z=-8 x+3y-2z=9 3x+2y+z=13

OpenStudy (candyme):

2x-4y+3z=-8 x+3y-2z=9 3x+2y+z=13 Correct?

OpenStudy (anonymous):

correct!!

OpenStudy (anonymous):

what do you need help with if you said it was correct?

OpenStudy (anonymous):

that was the question i need to solve the systems

OpenStudy (anonymous):

anyways you can use Gaussian elimination to solve this x + y + z = k x + y + z = k x + y + z = k you want it to look like this (k = constant value ) x + y + z = k 0 + y + z = k 0 + 0 + z = k example: just a name i am giving it | V R1: [ 2 x + 6 y + 1z = 1 ] R2 : 4x + 4y + 2z = 2 R3 : [ 2 x + 2y + 3z = 3 ] I want to use first (R)ow to cancel out multiply it with -1 i want to eliminate the 2 in R3 -1(R1) + R3 = R1: 2 x + 6 y + 1z = 1 R2 : 4x + 4y + 2z = 2 R3 : 0 x + 2y + 3z = 3 ------------------------------ you want to put it in this form so you can use the z substitute and solve x + y + z = 1 0 + y + z = 2 0 + 0 + z = 3

OpenStudy (anonymous):

wait what? im so confused

OpenStudy (anonymous):

my mom is even confused

OpenStudy (candyme):

I hope this helps you! Let me know if you have any other problems like this, I would be glad to help.